# Benford behavior resulting from stick and box fragmentation processes

### v2

@@ 3.4. Upper bound on error term $\mathcal{E}_{N}(a,b)$
$D_{2}(y)\ \ll\ N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}.$ (3.32)
Thus,
$\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)$ $\ \ll\ D_{1}(y)+D_{2}(y)\ \ll\ N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}+N^{-1/2-\delta}.$ (3.33)
We are now ready to bound the error term $\mathcal{E}_{N}(a,b)$ from (3.22):
$\mathcal{E}_{N}(a,b)$ $\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)dy$
$\ \ll\ N^{-\delta}\int_{-dC\sqrt{N}}^{dC\sqrt{N}}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}dy+\int_{-dC\cdot\sqrt{N}}^{dC\cdot\sqrt{N}}N^{-1/2-\delta}dy\ \ll\ N^{-\delta}.$ (3.34)
3.5. Strategy for remainder of proof
Our next task is to show that the main term $\mathcal{M}_{N}(a,b)$ defined in (3.21) satisfies
$\mathcal{M}_{N}(a,b)\ :=\ \sum_{-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy\ \approx\ b-a.$ (3.35)
Our strategy is the following. First, we want to write the integral of $d\mathcal{M}_{Y^{(N)}}(y)/dy$ over each interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ as the sum of a main term $m_{n}(a,b)$ that is constant with respect to $y$ within the interval and an error term $e_{n}(a,b)$ for each $n$. Essentially, this shows that $d\mathcal{M}_{Y^{(N)}}(y)/dy$ is almost constant within a small interval, incurring an negligible cost that, as we shall prove, does not accumulate, i.e.,
$\mathcal{M}_{\textup{err},N}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}e_{n}(a,b)\ \approx\ 0.$ (3.36)
@@ 3.5. Strategy for remainder of proof
$\mathcal{M}_{\textup{main},N}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}m_{n}(a,b)$ (3.37)
@@
$\mathcal{M}_{N}(a,b)\ =\ \mathcal{M}_{\textup{main},N}(a,b)+\mathcal{M}_{\textup{err},N}(a,b)\ \approx\ b-a.$ (3.38)
Thus, based on the estimate on the error term in (3.22),
$F_{N}(a,b)\ =\ \mathcal{M}_{N}(a,b)+\mathcal{E}_{N}(a,b)\ \approx\ b-a,$ (3.39)
@@
3.6. Equidistribution within small interval
In this section, we show that $d\mathcal{M}_{Y^{(N)}}(y)/dy$ is equidistributed within each small interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$. For each $n$, we write the integral over $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ in terms of a main term $m_{n}(a,b)$ that is constant with respect to $y$ and an error term $e_{n}(a,b)$, i.e.,
$\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy\ =\ m_{n}(a,b)+e_{n}(a,b),$ (3.40)
@@ 3.6. Equidistribution within small interval
$m_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)dy,$
$e_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)dy.$ (3.41)
Since $d\mathcal{M}_{Y^{(N)}}(n/\sqrt{N})/dy$ is constant, then the main term $m_{n}(a,b)$ is
$m_{n}(a,b)\ =\ \frac{b-a}{\sqrt{N}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right).$ (3.42)
@@
$\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)\right|,$ (3.43)
for $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]\supset[(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$. As we shall see, $d\mathcal{M}_{Y^{(N)}}(y)/dy$ is a continuous function over $n/\sqrt{N},(n+1)/\sqrt{N}]$ and a differentiable function over $(n/\sqrt{N},(n+1)/\sqrt{N})$. By the Mean Value Theorem, for any $y_{1},y_{2}\in[n/\sqrt{N},(n+1)/\sqrt{N}]$ with $y_{1}<y_{2}$, we have
$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{1})-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{2})\ =\ \frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(c_{n})(y_{1}-y_{2})$ (3.44)
for some $c_{n}\in(y_{1},y_{2})$. Hence
$\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{1})-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{2})\right|\ \leq\ \left|\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(c_{n})\right||y_{1}-y_{2}|\ \leq\ \frac{1}{\sqrt{N}}\left|\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(c_{n})\right|.$ (3.45)
We want to provide an upper on $|d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}|$ over the interval $n/\sqrt{N},(n+1)/\sqrt{N}]$. Recall from (3.2) that
@@
$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\quad\cdots dz_{m+1-d}$
@@
$\ =\ \ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (3.46)
Following the same procedure as shown in Section 3.2, we can again exchange the order of differentiation and integration. Since the product in the integrand in (3.6) goes from $j=2$ to d1d-1, we need to discuss the cases when $d=2$ and $d>2$ separately. If $d=2$,
$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)\ =\ \int_{-C\sqrt{N}}^{\frac{y}{2}}m_{Y^{{(N)}}}\left(z_{m-1},y-z_{m-1}\right)dz_{m-1}.$ (3.47)
By the Leibniz integral rule and (3.13),
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
$\ =\ \frac{1}{2}m_{Y^{(N)}}\left(\frac{y}{2},\frac{y}{2}\right)+\int_{-C\sqrt{N}}^{\frac{y}{2}}\frac{\partial}{\partial y}\left(m_{Y^{(N)}}(z_{m-1},y-z_{m-1})\right)dz_{m-1}$
@@
$\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-C\sqrt{N}}^{\frac{y}{2}}(\Phi(z_{m-1}))^{m-2}\varphi(z_{m-1})\cdot(-1)\cdot(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}.$ (3.48)
We know that $y$ is an upper bound on the value of $z_{m-1}+z_{m}$, then $y\geq-2C\sqrt{N}$, and $y/2\geq-C\sqrt{N}$. We now break into cases when $y\leq 0$ and $y\geq 0$. When $y\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi(y/2)$ on $C\sqrt{N},y/2]$. So (3.6) becomes
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|y-z_{m-1}|\varphi(y-z_{m-1})dz_{m-1}$
@@
$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1}\ \ll\ \varphi\left(\frac{y}{2}\right),$ (3.49)
@@
where in the last line we use the fact that the integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard normal random variable, which is finite. When $y\geq 0$, we have
$y-z_{m-1}$ $\ \geq\ y-\frac{y}{2}\ =\ \frac{y}{2}\ \geq\ 0.$ (3.50)
@@
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-C\sqrt{N}}^{\frac{y}{2}}(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}$
$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi(y-z_{m-1})\Bigg{|}_{z_{m-1}=-C\sqrt{N}}^{z_{m-1}=\frac{y}{2}}\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+C\sqrt{N}\right),$ (3.51)
where in the second line we employ the fact that $\int(y-x)\varphi(y-x)dx=\varphi(y-x)$. Thus, if $d=2$, regardless of whether $y\leq 0$ or $y\geq 0$, $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ is on the order of $\varphi(y/2)+\varphi(y+C\sqrt{N})$. This concludes the estimate of $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ for $d=2$.
If $d>2$, we repeatedly apply Leibniz integral rule as we have demonstrated in Subsection 3.4 to obtain
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
$\quad\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\Bigg{(}\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}m_{Y^{(N)}}\Bigg{(}z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\Bigg{)}$
$\quad\quad\quad\quad dz_{m-1}\Bigg{)}dz_{m-2}\cdots dz_{m+1-d}$
$\quad\ \ll\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
@@
$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg{)}dz_{m-2}\cdots dz_{m+1-d}.$ (3.52)
Now, similar to the case when $d=2$, we estimate the second integrand in (3.6) by breaking down into cases when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$ and $(y-\sum_{i=1}^{d-2}z_{m+i-d})\geq 0$. We leave the details to Appendix C. Hence, we have $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ in (3.6) is bounded above by

### v3

@@ 3.4 Upper bound on error term $\mathcal{E}_{N}(a,b)$
$D_{2}(y)\ \ll\ N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}.$ (112)
Thus, combining the estimates (110), (B) on $D_{1}(y)$, the estimate (112) on $D_{2}(y)$, as well as (106), we have
$\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)$ $\ \ll\ D_{1}(y)+D_{2}(y)+N^{-2\delta^{\prime}}\ \ll\ N^{-1/2-\delta}+N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}+N^{-2\delta^{\prime}}.$ (113)
We are now ready to bound the error term $\mathcal{E}_{N}(a,b)$ from (101):
$\mathcal{E}_{N}(a,b)$ $\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)dy$
$\ \ll\ N^{-\delta}\int_{-dN^{\delta^{\prime}}}^{dN^{\delta^{\prime}}}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}dy+\int_{-dN^{\delta^{\prime}}}^{dN^{\delta^{\prime}}}N^{-1/2-\delta}+N^{-2\delta^{\prime}}dy\ \ll\ N^{-\delta^{\prime}},$ (114)
where we use the fact that $0<\delta^{\prime}<\delta<1/2$.
3.5 Strategy for remainder of proof
Our next task is to show that the main term $\mathcal{M}_{N}(a,b)$ defined in (100) satisfies
$\mathcal{M}_{N}(a,b)\ :=\ \sum_{-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy\ \approx\ b-a.$ (115)
Our strategy is the following. First, we want to write the integral of $dF_{Y^{(N)},\textup{ main}}(y)/dy$ over each interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ as the sum of a constant main term $m_{n}(a,b)$ together with an error term $e_{n}(a,b)$ which, as we shall prove, does not accumulate, i.e.,
$\mathcal{M}_{N,\textup{ error}}(a,b)\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}e_{n}(a,b)\ \approx\ 0.$ (116)
@@ 3.5 Strategy for remainder of proof
$\mathcal{M}_{N,\textup{ main}}(a,b)\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}m_{n}(a,b)$ (117)
@@
$\mathcal{M}_{N}(a,b)\ =\ \mathcal{M}_{N,\textup{ main}}(a,b)+\mathcal{M}_{N,\textup{ error}}(a,b)\ \approx\ b-a.$ (118)
Thus, based on the estimate of the error term in (101),
$F_{N}(a,b)\ =\ \mathcal{M}_{N}(a,b)+\mathcal{E}_{N}(a,b)\ \approx\ b-a,$ (119)
@@
3.6 Equidistribution within small interval
In this section, we show that $dF_{Y^{(N)},\textup{ main}}(y)/dy$ is equidistributed within each small interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$. For each $n$, we write the integral over $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ as the sum of a constant main term $m_{n}(a,b)$ and an error term $e_{n}(a,b)$, i.e.,
$\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy\ =\ m_{n}(a,b)+e_{n}(a,b),$ (120)
@@ 3.6 Equidistribution within small interval
$m_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)dy,$
$e_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)dy.$ (121)
Since $dF_{Y^{(N)},\textup{ main}}(n/\sqrt{N})/dy$ is constant, then the main term $m_{n}(a,b)$ is
$m_{n}(a,b)\ =\ \frac{b-a}{\sqrt{N}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right).$ (122)
@@
$\left|\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)\right|,$ (123)
for $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]\supset[(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$. Recall from (3.2) that
$\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)$
@@
$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}})$
@@
$\quad\ =\ \ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})$
@@
$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}}).$ (124)
We let
$I_{N}(y)$
$\quad\ :=\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (125)
We see that $I_{N}(y)$ is a continuous function over $n/\sqrt{N},(n+1)/\sqrt{N}]$ and a differentiable function over $(n/\sqrt{N},(n+1)/\sqrt{N})$. By the Mean Value Theorem, for any $y_{1},y_{2}\in[n/\sqrt{N},(n+1)/\sqrt{N}]$ with $y_{1}<y_{2}$, we have
$I_{N}(y_{1})-I_{N}(y_{2})\ =\ \frac{d}{dy}I_{N}(c_{n})(y_{1}-y_{2})$ (126)
for some $c_{n}\in(y_{1},y_{2})$. Hence,
$\left|\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y_{1})-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y_{2})\right|$ $\ \leq\ \left|I_{N}(y_{1})-I_{N}(y_{2})\right|+O(N^{-2\delta^{\prime}})$
$\ \leq\ \left|\frac{d}{dy}I_{N}(c_{n})\right||y_{1}-y_{2}|+O(N^{-2\delta^{\prime}})$
$\ \ll\ \frac{1}{\sqrt{N}}\left|\frac{d}{dy}I_{N}(c_{n})\right|+O(N^{-2\delta^{\prime}}).$ (127)
To obtain an estimate on $dI_{N}(y)/dy$, we differentiate $I_{N}(y)$ by following the same procedure as shown in Section 3.2 to exchange the order of differentiation and integration. Since the product in the integrand in (3.6) goes from $j=2$ to d1d-1, we need to discuss the cases when $d=2$ and $d>2$ separately. If $d=2$,
$I_{N}(y)\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{2}}f_{Z^{(N)}_{(m-1)},Z^{(N)}_{(m)},\textup{ main}}\left(z_{m-1},y-z_{m-1}\right)dz_{m-1}.$ (128)
By the Leibniz integral rule and (81),
$\frac{d}{dy}I_{N}(y)$
$\quad\ =\ \frac{1}{2}f_{Z^{(N)}_{(m-1)},Z^{(N)}_{(m)},\textup{ main}}\left(\frac{y}{2},\frac{y}{2}\right)+\int_{-N^{\delta^{\prime}}}^{\frac{y}{2}}\frac{\partial}{\partial y}\left(f_{Z^{(N)}_{(m-1)},Z^{(N)}_{(m)},\textup{ main}}(z_{m-1},y-z_{m-1})\right)dz_{m-1}$
@@
$\quad\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-N^{\delta^{\prime}}}^{\frac{y}{2}}(\Phi(z_{m-1}))^{m-2}\varphi(z_{m-1})\cdot(-1)\cdot(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}.$ (129)
We know that $y\geq-2N^{\delta^{\prime}}$, so $y/2\geq-N^{\delta^{\prime}}$. We now break into cases when $y\leq 0$ and $y\geq 0$. When $y\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi(y/2)$ on $N^{\delta^{\prime}},y/2]$. So (3.6) becomes
$\frac{d}{dy}I_{N}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|y-z_{m-1}|\varphi(y-z_{m-1})dz_{m-1}$
$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1}\ \ll\ \varphi\left(\frac{y}{2}\right),$ (130)
@@
where in the last line we use the fact that the integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard Gaussian random variable, which is finite. When $y\geq 0$, we have
$y-z_{m-1}$ $\ \geq\ y-\frac{y}{2}\ =\ \frac{y}{2}\ \geq\ 0.$ (131)
@@
$\frac{d}{dy}I_{N}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-N^{\delta^{\prime}}}^{\frac{y}{2}}(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}$
$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi(y-z_{m-1})\Bigg|_{z_{m-1}=-N^{\delta^{\prime}}}^{z_{m-1}=\frac{y}{2}}\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+N^{\delta^{\prime}}\right),$ (132)
where in the second line we employ the fact that $\int(y-x)\varphi(y-x)dx=\varphi(y-x)$. Thus, if $d=2$, regardless of whether $y\leq 0$ or $y\geq 0$, $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ is on the order of $\varphi(y/2)+\varphi(y+N^{\delta^{\prime}})$. This concludes the estimate of $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ for $d=2$.
If $d>2$, we repeatedly apply Leibniz integral rule as we have demonstrated in Subsection 3.4 to (3.6) and also use (81) to obtain
$\frac{d}{dy}I_{N}(y)$
$\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\Bigg(\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\Bigg(z_{m+1-d},\dots,z_{m-1},$
$\quad\quad y-\sum_{i=1}^{d-1}z_{m+i-d}\Bigg)dz_{m-1}\Bigg)dz_{m-2}\cdots dz_{m+1-d}$
$\ \ll\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
@@
$\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg)dz_{m-2}\cdots dz_{m+1-d}.$ (133)
Now, similar to the case when $d=2$, we estimate the second integrand in (3.6) by breaking down into cases when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$ and $(y-\sum_{i=1}^{d-2}z_{m+i-d})\geq 0$. We leave the details to C. Hence, we have $dI_{N}(y)/dy$ in (3.6) is bounded above by